A partical is moving in a circle of radious R in such a way that at any instant the total acceleration makes an angle of 450 with radious.initial speed of partical is V0.the time taken to complete the first revolution is-
  1. (R/V0 e-2pi​),
  2. R/V(1-e-2pi)
  3. R/V0
  4. 2R/V0

Grade:12th pass

1 Answers

Vikas TU
6 years ago
Dear Student,
As we know that,
v2/R=acos
=>V02/R = αR/√2
=>√2V02/R2 = α
=αt2/2
=>2π=√2V02/R2 t2
=>t=√(2X√2 πR^2)/V02)
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)